# Static and Dynamic BART for Rank-Order Data

### v4

@@ 5 Real Data Applications
The usefulness of our static data models is demonstrated by an application to three standard label ranking benchmark datasets (de Sá et al., 2018). We include in these results a variation on ROBART, named SoftROBART, with soft and sparse splitting rules (Linero and Yang, 2018). Then, we compare static and dynamic model specifications on the two time series datasets presented in Section 1, about country-rankings on the economic complexity index and the 2022 NCAA rankings.
@@ 5.1 Static Label Rankings
The sushi dataset includes 5000 rankings of 10 items: shrimp, sea eel, tuna, squid, sea urchin, salmon roe, egg fatty tuna, tuna roll and cucumber roll (Kamishima et al., 2010). The variables include age, gender, answer duration, a binary variable indicating if the respondent moved to another city, and categorical variables for current and past region and prefecture. The categorical variables were re-coded as dummy variables resulting in 128 predictors.
The Top7Movies dataset is a subset of the MovieLens dataset (Harper and Konstan, 2016) processed by de Sá et al. (2018). The rankings were constructed from ratings of 7 movies submitted by 602 individuals. Predictors include gender, age, latitude and longitude of zipcode, city, state, and occupation. The categorical variables for city, state, and occupation were coded as dummy variables, resulting in a total of 446 predictors. This dataset contains many ties, and therefore allows for a comparison of methods in an application to partial ranking data.
@@ 5.2 Dynamic Label Ranking: Economic Complexity Index
The ECI dataset (Alvo and Philip, 2014; Liu et al., 2019) includes data for 80 countries from 1965 to 2015. The dependent variable is the country-year specific ranking in an index defined by the Observatory of Economic Complexity. The independent variables include the lag of the observed rank, GDP per capita, population, population growth, cereal production in tonnes, the dependency ratio, and growth in the dependency ratio. These are not necessarily the most relevant variables for the index of interest. However, this exercise demonstrates the ability of the new methods to select relevant variables.
Table 7 shows that standard ROBART and ARROBART do not improve on a static linear model. However, RODART and ARRODART, variations on ROBART and ARROBART for which a Dirichlet hyperprior distribution is specified for sparse splitting rules (Linero, 2018), improve substantially on the linear models. Specifically, the dynamic ARRODART model provides better forecasts than the static RODART model. We refer to the Supplement for the data sources and further results.
@@ 5.3 Panel of Label Rankings: NCAA Pollsters
To inspect in more detail the advantages of the BART specification compared to the linear one, we report the point forecasts from the best ARROBART and ARROLinear models in Fig. 2 against the observed ranks, for $N=7$ teams and $h=5$ test periods. As the rankers are assumed to be independent, we present the results for a single pollster. The ARROBART model is able to correctly predict the rank for several teams, with minor deviations for Alabama and Clemson. Conversely, the ARROLinear forecasts are generally more distant from the actual ranks. In particular, it can be observed that while both methods often produce similar forecasts, ARROBART produces more accurate forecasts for Alabama, USC, and Michigan in weeks 12 to 14.
The point forecast falls short of quantifying the uncertainty about the predicted value. To tackle this issue, we investigate in deeper detail the posterior predictive distribution of the best-performing model. Figure 3 reports the posterior predictive distribution of the rank for Alabama and Michigan, whose observed ranks have changed multiple times during the testing sample (see Fig. 2). For both teams, the predictive density of the rank is unimodal and changes over time in terms of its mode and dispersion, suggesting a horizon-specific uncertainty of predictions. Furthermore, the comparison of Fig. 2 and Fig. 3 highlights that in those periods where the point forecast differs from the observed one, the latter falls within the second-most frequent rank of the posterior predictive distribution.
@@ 6 Discussion
This article introduces a novel class of nonparametric Thurstone models for static and dynamic regressions for rank-order data called ROBART and ARROBART. The proposed frameworks use a BART specification for the regression function of the latent variable representation to allow for flexible, possibly nonlinear relationships between the latent scores and their lagged values as well as exogenous covariates. The filtering and smoothing recursions for the dynamic model are derived, followed by a discussion on the computational tools needed to sample from them. A simulation study illustrates that the proposed ROBART and ARROBART outperform the competitors in forecasting. Then, ROBART is applied to rankings of sushi, movies, and political parties, and ARROBART is applied to a time series of economic complexity and NCAA rankings. The results highlight the better performance of the ROBART and ARROBART models compared to linear and static counterparts, suggesting the importance of accounting for temporal persistence and the usefulness of BART in accounting for the nonlinearities driving the evolution of the latent item-specific scores.
The proposed methods can be applied to a wide range of datasets. Examples include the analysis of the main drivers of customers’ preferences, university rankings, and political ratings, in addition to sports data. The proposed ARROBART model can be generalized to $p>1$ lags in the presence of more complex dynamics of the latent scores. Other interesting possible extensions of the proposed methods include investigating the performance of more general BART specifications. For instance, the introduction of a varying-coefficient BART (Deshpande et al., 2020) or partially linear BART model for the latent scores (Zeldow et al., 2019). The proposed autoregressive BART methods can also be adapted to investigate binary, ordered categorical, unordered categorical data, or censored outcome data. These models involve an autoregressive Gaussian process for the latent scores and are therefore similar to the rank-order model described in this article. The samplers can be derived in a similar way to the sampler in this work, and have simpler constraints on the latent variables than rank-order outcome models. Finally, it would be interesting to extend the proposed dynamic model to investigate the problem of aggregation of ranking lists (Deng et al., 2014; Zhu et al., 2023) in a time series setting.

### v5

@@ 5 Real Data Applications
The usefulness of our static data models is demonstrated by an application to three standard label ranking benchmark datasets (de2018discovering). We include in these results a variation on ROBART, named SoftROBART, with soft and sparse splitting rules (linero2018dart). Then, we compare static and dynamic model specifications on the two time series datasets presented in Section 1, about country-rankings on the economic complexity index and the 2022 NCAA rankings.
@@ 5.1 Static Label Rankings
The sushi dataset includes 5000 rankings of 10 items: shrimp, sea eel, tuna, squid, sea urchin, salmon roe, egg fatty tuna, tuna roll and cucumber roll (kamishima2010survey). The variables include age, gender, answer duration, a binary variable indicating if the respondent moved to another city, and categorical variables for current and past region and prefecture. The categorical variables were re-coded as dummy variables resulting in 128 predictors.
The Top7Movies dataset is a subset of the MovieLens dataset (harper2016movielens) processed by de2018discovering. The rankings were constructed from ratings of 7 movies submitted by 602 individuals. Predictors include gender, age, latitude and longitude of zipcode, city, state, and occupation. The categorical variables for city, state, and occupation were coded as dummy variables, resulting in a total of 446 predictors. This dataset contains many ties, and therefore allows for a comparison of methods in an application to partial ranking data.
@@ 5.2 Dynamic Label Ranking: Economic Complexity Index
The ECI dataset (alvo2014statistical; liu2019model) includes data for 80 countries from 1965 to 2015. The dependent variable is the country-year specific ranking in an index defined by the Observatory of Economic Complexity. The independent variables include the lag of the observed rank, GDP per capita, population, population growth, cereal production in tonnes, the dependency ratio, and growth in the dependency ratio. These are not necessarily the most relevant variables for the index of interest. However, this exercise demonstrates the ability of the new methods to select relevant variables.
Table 7 shows that standard ROBART and ARROBART do not improve on a static linear model. However, RODART and ARRODART, variations on ROBART and ARROBART for which a Dirichlet hyperprior distribution is specified for sparse splitting rules (linero2018bayesian), improve substantially on the linear models. Specifically, the dynamic ARRODART model provides better forecasts than the static RODART model. We refer to the Supplement for the data sources and further results.
@@ 5.3 Panel of Label Rankings: NCAA Pollsters
To inspect in more detail the advantages of the BART specification compared to the linear one, we report the point forecasts from the best ARROBART and ARROLinear models in Fig. S.8 against the observed ranks, for $N=7$ teams and $h=5$ test periods. As the rankers are assumed to be independent, we present the results for a single pollster. The ARROBART model is able to correctly predict the rank for several teams, with minor deviations for Alabama and Clemson. Conversely, the ARROLinear forecasts are generally more distant from the actual ranks. In particular, it can be observed that while both methods often produce similar forecasts, ARROBART produces more accurate forecasts for Alabama, USC, and Michigan in weeks 12 to 14.
The point forecast falls short of quantifying the uncertainty about the predicted value. To tackle this issue, we investigate in deeper detail the posterior predictive distribution of the best-performing model. Figure S.9 reports the posterior predictive distribution of the rank for Alabama and Michigan, whose observed ranks have changed multiple times during the testing sample (see Fig. S.8). For both teams, the predictive density of the rank is unimodal and changes over time in terms of its mode and dispersion, suggesting a horizon-specific uncertainty of predictions. Furthermore, the comparison of Fig. S.8 and Fig. S.9 highlights that in those periods where the point forecast differs from the observed one, the latter falls within the second-most frequent rank of the posterior predictive distribution.
@@ 6 Discussion
This article introduces a novel class of nonparametric Thurstone models for static and dynamic regressions for rank-order data called ROBART and ARROBART. The proposed frameworks use a BART specification for the regression function of the latent variable representation to allow for flexible, possibly nonlinear relationships between the latent scores and their lagged values as well as exogenous covariates. The filtering and smoothing recursions for the dynamic model are derived, followed by a discussion of Gibbs sampling methods. A simulation study illustrates that the proposed ROBART and ARROBART outperform the competitors in forecasting. Then, ROBART is applied to rankings of sushi, movies, and political parties, and ARROBART is applied to a time series of economic complexity and NCAA rankings. The results highlight the better performance of the ROBART and ARROBART models compared to linear and static counterparts, suggesting the importance of accounting for temporal persistence and the usefulness of BART in accounting for the nonlinearities driving the evolution of the latent item-specific scores.
The proposed methods can be applied to a wide range of datasets. Examples include the analysis of the main drivers of customers’ preferences, university rankings, and political ratings, in addition to sports data. Based on the results from both simulations and real data applications, we encourage the use of the ROBART model or its variants to investigate static (full or partial) ranking data. Conversely, the ARROBART model or its variants are applicable to dynamic ranking data. Moreover, if the data shows sparsity, (AR)RODART should be used, and in the absence of item-specific covariates, (AR)ROBART/(AR)RODART with item-specific trees shall be preferred.
The proposed ARROBART model can be generalized to $p>1$ lags in the presence of more complex dynamics of the latent scores. Other interesting possible extensions of the proposed methods include investigating the performance of more general BART specifications. For instance, the introduction of a varying-coefficient BART (deshpande2020vcbart) or partially linear BART model for the latent scores (zeldow2019semiparametric). The proposed autoregressive BART methods can also be adapted to investigate binary, ordered categorical, unordered categorical data, or censored outcome data. These models involve an autoregressive Gaussian process for the latent scores and are therefore similar to the rank-order model described in this article. The samplers can be derived in a similar way to the sampler in this work, and have simpler constraints on the latent variables than rank-order outcome models. Finally, it would be interesting to extend the proposed dynamic model to investigate the problem of aggregation of ranking lists (deng2014bayesian; zhu2023partition) in a time series setting.